The reaction time of a driver to visual stimulus is normally distributed with a mean of 0.4

seconds and a standard deviation of 0.05 seconds. Use R to find the

(a) probability that a reaction requires more than 0.5 seconds.

(b) probability that a reaction requires between 0.4 and 0.5 seconds.


1
Expert's answer
2022-04-20T13:40:20-0400

Denote by "X" the random variable that has a normal distribution with parameters "\mu=0.4" and "\sigma=0.05". The probability density function is: "f(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}" .

(a) "P(X\geq0.5)=\int_{0.5}^{+\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}dx\approx0.0228"

(b) "P(0.4\leq X\leq0.5)=\int_{0.4}^{0.5}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}dx\approx0.4772"

The answers are rounded to 4 decimal places